Simplify and write the answer in scientific notation. 0.0885 : 1,770,000 x 10

Topics of contemporary mathematics
Answer #1

Solution 0.0885=1770,000 → 885*10* = 1770x103 1885X104 + 2x888X109 1x10+ = 2X103 loxios = 24103 - 5x105 = 1x103 -5 - 3 7 5X10